OpenStudy (waheguru):

Please help me How to find the recursive formula for 5,10, 20, 40...

OpenStudy (waheguru):

The general equation is t(n) = t(n-1)+d

jimthompson5910 (jim_thompson5910):

you're using the wrong recursive general equation you should be using this t(n) = r*t(n-1) where t(n) is the nth term, r is the common ratio, and t(n-1) is the term before the nth term

OpenStudy (waheguru):

Would this be correct then? t(n) = t(n-1)*2 ?

jimthompson5910 (jim_thompson5910):

or t(n) = 2*t(n-1) good

OpenStudy (waheguru):

Thanks and what about the general term formula?

jimthompson5910 (jim_thompson5910):

well this is a geometric sequence with first term a = 5, common ratio r = 2

jimthompson5910 (jim_thompson5910):

so you tell me

OpenStudy (waheguru):

t(n) = 5 + something?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

the nth term of a geometric sequence is found by t(n) = a*r^(n-1)

OpenStudy (waheguru):

What does r stand for?

jimthompson5910 (jim_thompson5910):

r is the common ratio

jimthompson5910 (jim_thompson5910):

to find the next term, you multiply by the common ratio

OpenStudy (waheguru):

Ok, so I got t = 5*2(n-1)

OpenStudy (waheguru):

I understand this know, i appreciate your help alot

jimthompson5910 (jim_thompson5910):

more like \[\Large t(n) = 5*2^{n-1}\] but I think that's what you meant to say

OpenStudy (waheguru):

Yes, thanks

jimthompson5910 (jim_thompson5910):

yw